OpenStudy (anonymous):

Why is this answer wrong:

OpenStudy (anonymous):

Here are my steps and my answer:

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin ^{5}xdx\]

OpenStudy (anonymous):

= \[\int\limits_{}^{}\sin x(1-\cos ^{4}x)\]

OpenStudy (anonymous):

U sub: u = cosx du=-sinxdx -> -du=sinxdx

OpenStudy (anonymous):

intsinx=-cosx then -sinxcos^4x=1/5cos^5x so all together you need 1/5cos^5x-cosx+C

OpenStudy (anonymous):

\[-\int\limits_{}^{}1-u ^{4}\] = \[-u+u ^{5}/5\]

OpenStudy (anonymous):

So my final answer would be:

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[−cosx+\cos5x/5 + c\]

OpenStudy (anonymous):

But the answer give on pauls notes is not the same?

OpenStudy (anonymous):

what does it say on there?

OpenStudy (anonymous):

So is my methodology incorrect or did it just yield a different form of the correct answer

OpenStudy (anonymous):

example 1

OpenStudy (anonymous):

he did a different trig. sub.

OpenStudy (anonymous):

ok do you see where he does sin^5x=(1-u^2)^2 is wrong he is missing something cause that would yeild sin^4x

OpenStudy (anonymous):

He has the other sinx include outside of that parenthesis

OpenStudy (anonymous):

If you look before the u sub

OpenStudy (anonymous):

sin^4x does not equal 1-cos^4x thats what you did wrong he is rightt

OpenStudy (anonymous):

Why not? isnt that the fundamental trig.

OpenStudy (anonymous):

ohhh yea

OpenStudy (anonymous):

it only works for sin^2x+cos^2x

OpenStudy (anonymous):

1-2cos^2x_cos^4x

OpenStudy (anonymous):

Gotta remember that, thanks mate

OpenStudy (anonymous):

so its easy once you do that but THE INTEGRAL YOU DID BEFORE THIS IS CORRECT KEEP WORKING HARD THESE ARE FAILY SIMPLE ONCE YOU GET THE HANG OF THEM